18,000-100P1 and the demand function for the second group are given as Q2 = 17,000-100P2. The marginal cost for each group is $5.00. What will be the profit, if they cannot successfully segment the markets for the two groups of people? What will be the profit if they can successfully discriminate against 2 groups of people?


1
Expert's answer
2021-04-08T05:28:29-0400
"Q_1=18,000-100p_1"

"Q_2=17,000-100p_2"

"Q=35,000-100p_1-100p_2"

"MC=p_1=p_2=5"

"Q=35,000-200\times 5=34,000"

"MC=p_1=5"

"Q=18,000-100\times5=17,500"